OpenStudy (anonymous):

1/3(y+3)=4y: i got 11/3 the answer is 3/11 how do i flip it?

OpenStudy (anonymous):

1/3(y+3)=4y (y+3)/3=4y y+3=12y 3=11y y=3/11

OpenStudy (anonymous):

the term (y+3) is in numerator or denominator or numerator?

OpenStudy (anonymous):

it is 1/3 mutiplied by (y+3)=4y

OpenStudy (lalaly):

is it\[\frac{1}{3} (y+3)=4y\]

OpenStudy (anonymous):

yes

OpenStudy (lalaly):

\[\frac{y+3}{3}=4y\]now multiply both sides by 3\[\cancel{3} \times \frac{y+3}{\cancel{3}} =4y \times 3\]\[y+3=12y\]\[11y=3\]\[y=\frac{3}{11}\]

OpenStudy (lalaly):

you dont have to flip anything